OpenStudy (pattylove22):

In the week before and the week after a holiday, there are 11,000 total deaths, and 5467 of them occurred in the week before the holiday. Construct a 90% confidence interval estimate of the proportion of deaths in the week before the holiday to the total deaths in the week before and the week after the holiday.

OpenStudy (loviy):

Take note of the formula: \[\displaystyle \hat{p} \pm \text{z}_{\alpha/2} * (\sqrt{\displaystyle \frac{\hat{p} * (1-\hat{p})}{n}})\] where p-hat is the sample proportion. \[\hat p = \frac{x}{n}\] where x is the number of successes and n is the sample size